# Design and development of optical modules for the BUTTON-30 detector

### v1

@@ 2 Design of the optical module
High-voltage (HV) and signal return from the PMT are provided by a 25 m coaxial cable integrated into the PMT base. This exits the housing via a watertight assembly piece refered to as the penetrator system. This consists of a body piece, two Viton O-rings, a ferrule, and two nuts. Tightening the outer nut compresses the ferrule against one of the O-rings, ensuring a watertight seal around the cable. A silica desiccator bag attached to the PMT base controls the humidity level within the housing. The inner surface of the rear of the housing is painted black with acrylic paint to block the reflection of stray-light. The chosen paint, RS Matt-Black Aerosol, reduces light transmission for wavelengths near $400\,\textrm{nm}$ by $99\,\%$.
@@ 4 Pressure qualification
The first prototypes were produced by vacuum thermoforming of acrylic. This technique resulted in variations in thickness and flatness, particularly in the transition region between the dome and the flange, where stresses are concentrated. Under 3 bar applied pressure, the three-dimensional COMSOL model predicts a maximum stress of $\sim$60 MPa in this region due to abrupt thinning (Figure 3). Three prototype modules were tested in a high-pressure vessel. All were able to sustain pressures up to 3 bar. One module was kept at 2 bar for 12 hours without signs of water ingress (verified using humidity paper placed inside the housing). Destructive implosion of one module was observed at 3.2 bar overpressure, consistent with expectations.
Although the pressure strength of the thermoformed housings exceeds the requirements of BUTTON-30 by an order of magnitude, the prototype housings were fragile. In particular, the flange was susceptible to cracking. To address this, ICLTech switched to a blow-moulding process for the final production. This change resulted in a thicker flange region with greater uniformity, significantly improving the durability and handling characteristics. The wall thickness for the blow-moulded design gradually decreases to approximately 33\,$m$ near the apex of the hemisphere. Despite this, the COMSOL simulation (Figure 4) shows no significant increase in stress at the apex. In addition, the maximum stress around the flange is reduced to $\sim$20 MPa as seen in Figure 4. The measured pressure strength of the prototypes and the improved robustness to handling of the blow-moulded housings give us confidence that the final housings meet the needs of BUTTON-30.
@@ 5.2 Degassing
Once degassing was complete, the acrylic lid was removed. The PMT, secured in the holder, was lowered into the gel. The rig and holder were adjusted so that a gap of 4 mm was maintained between the photocathode and the front hemisphere, allowing the gel to spread uniformly over the photocathode surface. This procedure was completed within the 30-minute work time of the mixed gel specified by the manufacturer. Once positioned, the PMT and holder were left for 24 hours to allow the gel to cure.
@@ 5.4 Torquing
The bolts were tightened in stages: an initial pass at 4.5 Nm, followed by two passes at 6 Nm, using a cross pattern to ensure uniform tension across the flange. Finally, the outer cap nut of the penetrator was tightened to 12 Nm. A checklist was maintained to confirm that this procedure was followed. A fully assembled module is shown in Figure 8.
@@ 5.5 Quality assurance and shipping
As a further check, a spare module was shipped to Brookhaven National Laboratory. This module was operated in WbLS for several weeks with no visible signs of corrosion or degradation.

### v3

@@ 2 Design of the optical module
High-voltage (HV) and signal return from the PMT are provided by a 25 m coaxial cable integrated into the PMT base. The cable exits the housing via a watertight assembly referred to as the penetrator system. This consists of a body piece, two Viton O-rings, a ferrule, and two nuts. Tightening the outer nut compresses the ferrule against one of the O-rings, ensuring a watertight seal around the cable. A silica desiccator bag attached to the PMT base controls the humidity level within the housing. The inner surface of the rear of the housing is painted black with acrylic paint to block the reflection of stray-light. The chosen paint, RS Matt-Black Aerosol, reduces light transmission for wavelengths near $400\,\textrm{nm}$ by $99\,\%$.
@@ 4 Pressure qualification
The first prototypes were produced by vacuum thermoforming of acrylic. This technique resulted in variations in thickness and flatness, particularly in the transition region between the dome and the flange, where stresses are concentrated. Under 3 bar applied pressure, the three-dimensional COMSOL model predicts a maximum stress of $\sim$60 MPa in this region due to abrupt thinning (Figure 3). Three prototype thermoformed modules were tested in a high-pressure vessel. All were able to sustain pressures up to 3 bar. One module was kept at 2 bar for 12 hours without signs of water ingress (verified using humidity paper placed inside the housing). Destructive implosion of one thermoformed module was observed at 3.2 bar overpressure, consistent with expectations.
Although the pressure strength of the thermoformed housings exceeds the requirements of BUTTON-30 by an order of magnitude, the prototype housings were fragile. In particular, the flange was susceptible to cracking during handling, which resulted in a low production yield. To address this, ICLTech switched to a blow-moulding process for the final production. This change resulted in a thicker flange region with greater uniformity, significantly improving the durability and handling characteristics. The wall thickness for the blow-moulded design gradually decreases to approximately 33\,$m$ near the apex of the hemisphere. Despite this, the COMSOL simulation (Figure 4) shows no significant increase in stress at the apex. In addition, the maximum stress around the flange is reduced to $\sim$20 MPa as seen in Figure 4. Due to time constraints and the unavailability of the pressure test chamber, destructive tests on the new blow-moulded housings could not be carried out. Nevertheless, based on the performance of the previous iteration and the improved handling characteristics, we are confident that the new housings meet BUTTON’s requirements.
@@ 5.2 Degassing
Once degassing was complete, the acrylic lid was removed. The PMT, secured in the holder, was lowered into the gel. The rig and holder were adjusted so that a gap of 4 mm was maintained between the photocathode and the front hemisphere, allowing the gel to spread uniformly over the photocathode surface. This procedure was completed within the 30-minute work time of the mixed gel specified by the manufacturer. Once positioned, the PMT and holder were left for 24 hours to allow the gel to cure. The room temperature and humidity during curing was around $24\pm 2~^{\circ}\mathrm{C}$ and $45\pm 5\%$ respectively.
@@ 5.4 Torquing
The bolts were tightened in stages using torque wrenches ($4\%$ tolerance): an initial pass at 4.5 Nm, followed by two passes at 6 Nm, using a cross pattern to ensure uniform tension across the flange. Finally, the outer cap nut of the penetrator was tightened to 12 Nm. A checklist was maintained to confirm that this procedure was followed. A fully assembled module is shown in Figure 8.
@@ 5.5 Quality assurance and shipping
As a further check, a spare module was shipped to Brookhaven National Laboratory (BNL). This module was operated in $1\%$ WbLS for several weeks with no visible signs of corrosion, degradation or cracks.